OpenStudy (anonymous):

how do you rationalize the fraction: 1/(5 x √5−3 y √3) please.

OpenStudy (diyadiya):

multiply denominator by \[5x \sqrt{5}+3y \sqrt{3} \] and numerator by \[5x \sqrt{5} -3y \sqrt{3} \]

OpenStudy (anonymous):

I did that and i got the numerator of course, what would you get for the denomiator, i think that is what's confusing me im not sure if anything cancels out

OpenStudy (diyadiya):

(a-b)(a+b) = a^2 - b^2 so numerator will be 125x^2 - 27 y^2

OpenStudy (anonymous):

125x^2 - 27y^2

OpenStudy (diyadiya):

i meant denominator

OpenStudy (diyadiya):

(5x√5)^2=25*5x^2=125x^2 (3y√3)^2=9*3x^2=27y^2

OpenStudy (anonymous):

thank you very much and the numerator would just be 5x5√+3y3√ then correct?

OpenStudy (diyadiya):

yes :)

OpenStudy (anonymous):

alrighty thank you!

OpenStudy (diyadiya):

yw:)